15.8 g of KCl is dissolved in 225 mL of water. Calculate the molarity.
1
Expert's answer
2018-02-21T04:50:41-0500
molarity = moles of solute/liters of solution
M(KC) = 39+35.5 = 74.5 g/mol
n(KCl) = 15.8 g / 74.5 g/mol = 0.212 moles
V(KCl) = n*Vm = 0.212 moles*22.4L/mol = 4.75 L
C = 0.212moles / (0.225+4.75)L = 0.212 moles / 4.975 L = 0.0426 mol/L